The base of a triangle is one more than four times the height. Determine the dimensions that will give a total area of 9 cm^2. What is the minimum area of such a triangle?
Answer by RAY100(1637)

Let h= height of triangle
let b=base= (4h+1)
.
Area = 1/2 *b*h = 1/2 * (4h+1)*h = 2h^2 + 1/2 h
.
but Area = 9 = 2h^2 +h/2
.
or,,,,0= 2h^2 +h/2 -9,,,,,,,mult by 2
.
0= 4h^2 +h - 18
.
0 = (4h+9)(h-2)
.
h=2,,,,-9/4{not realistic}
.
h=2,,,b=9
.
.
check a=1/2 * 2*9 =9,,,,,ok
b=4*2+1=9,,,,,,ok